OpenStudy (anonymous):

Solve the following equation: a-4/2(a-1)-3/5=0

OpenStudy (anonymous):

a= -17

OpenStudy (anonymous):

a-4/2a-2 = 3/5 a-4 = 3/5 x (2a-2) a-4 = (6a-6)/5 5(a-4) = 6a-6 5a-20 = 6a-6 -20+6 = 6a-5a -14 = a check -14-4/2(-14-1)-3/5 = -18/-30 -3/5 = 0 note: -18/-30 becomes 3/5 when divided by -6/-6 ( since -6/-6 cancels out the fraction to its lowest terms) answer : a = -14